# Rethinking Discrete Tokens: Treating Them as Conditions for Continuous Autoregressive Image Synthesis

### v1

@@ Abstract
Recent advances in large language models (LLMs) have spurred interests in encoding images as discrete tokens and leveraging autoregressive (AR) frameworks for visual generation. However, the quantization process in AR-based visual generation models inherently introduces information loss that degrades image fidelity. To mitigate this limitation, recent studies have explored to autoregressively predict continuous tokens. Unlike discrete tokens that reside in a structured and bounded space, continuous representations exist in an unbounded, high-dimensional space, making density estimation more challenging and increasing the risk of generating out-of-distribution artifacts. Based on the above findings, this work introduces DisCon (Discrete-Conditioned Continuous Autoregressive Model), a novel framework that reinterprets discrete tokens as conditional signals rather than generation targets. By modeling the conditional probability of continuous representations conditioned on discrete tokens, DisCon circumvents the optimization challenges of continuous token modeling while avoiding the information loss caused by quantization. DisCon achieves a gFID score of 1.38 on ImageNet 256$\times$256 generation, outperforming state-of-the-art autoregressive approaches by a clear margin.
@@ 2.2 Discrete Autoregressive Image Generation
Inspired by the success of large language models [21, 36, 33, 1, 18] in modeling discrete sequences, autoregressive (AR) approaches have been adapted to image generation [44, 15, 22, 11, 10, 52, 47, 38, 28, 45, 30]. These methods quantize images into sequences of discrete tokens (e.g., via VQGAN [44]) and model the token distribution sequentially using cross-entropy loss. While this enables efficient training and fast inference, the quantization process inevitably discards fine-grained image details, limiting reconstruction fidelity. Notable examples include RAR [46], which refines generation by permuting token sequences for richer context, and VAR [34], which models image structure through next-token prediction. However, the inherent limitations of discrete tokens restrict the expressive capacity needed for high-fidelity synthesis.
@@ 3.3 Proposed Method: DisCon
$\hat{\mathbf{x}}_{d}=\text{DiscreteAR}(\mathbf{I}).$ (6)
@@
Why DisCon Works. We decompose the modeling of continuous tokens into two subproblems. First, we model the discrete tokens $\mathbf{x}_{d}$ using a discrete AR model trained with a cross-entropy loss, a formulation that has been proven effective in prior work. Second, we model the conditional probability from discrete tokens to continuous tokens, i.e., $p(\mathbf{x}_{c}|\mathbf{x}_{d})$, using a continuout AR model. Formally, we factorize the overall distribution as
@@ 4.1 Implementation Details
Table 1: Quantitative comparisons on ImageNet-256. Our method achieves SOTA performance among AR models. Among the metrics, gFID is the most important metric for evaluating the fidelity of the synthesis result. Params denotes the number of trainable parameters.
@@ 4.3 Ablation Studies
Table 2: Ablation study results on discrete token conditioning. We condition MAR on discrete tokens across various settings, including different parameter sizes and AR steps. Note that these results are preliminary; additional training time and AR steps are expected to further improve performance. The sec/image values were measured using a batch size of 100, and our method’s inference time accounts for the discrete token generation.
@@
Figure 8: Generation Performance during Training. Our method achieves SOTA performance (around 1.4 gFID) after 200 training epochs. Note that these results are obtained with a 0.9999 EMA setting, which may cause a slight performance lag. Detailed training loss curves in the supplementary materials provide additional insights.
@@
Figure 8 shows the generation performance of our method during training. Our model achieves SOTA performance (around 1.4 gFID) after only 200 training epochs, although the use of a 0.9999 EMA (Exponential Moving Average) may introduce a performance lag. This training efficiency demonstrates the reduced optimization complexity achieved by modeling the conditional probability from the discrete distribution to the continuous distribution. Detailed training loss curves in the supplementary materials further support these findings.

### v2

@@ Abstract
Recent advances in large language models (LLMs) have spurred interests in encoding images as discrete tokens and leveraging autoregressive (AR) frameworks for visual generation. However, the quantization process in AR-based visual generation models inherently introduces information loss that degrades image fidelity. To mitigate this limitation, recent studies have explored to autoregressively predict continuous tokens. Unlike discrete tokens that reside in a structured and bounded space, continuous representations exist in an unbounded, high-dimensional space, making density estimation more challenging and increasing the risk of generating out-of-distribution artifacts. Based on the above findings, this work introduces DisCon (Discrete-Conditioned Continuous Autoregressive Model), a novel framework that reinterprets discrete tokens as conditional signals rather than generation targets. By modeling the conditional probability of continuous representations conditioned on discrete tokens, DisCon circumvents the optimization challenges of continuous token modeling while avoiding the information loss caused by quantization. DisCon achieves a gFID score of 1.38 on ImageNet 256$\times$256 generation, outperforming state-of-the-art autoregressive approaches by a clear margin. Project page: https://pengzheng0707.github.io/DisCon.
@@ 2.2 Discrete Autoregressive Image Generation
Inspired by the success of large language models [21, 36, 33, 1, 18] in modeling discrete sequences, autoregressive (AR) approaches have been adapted to image generation [44, 15, 22, 11, 10, 52, 47, 38, 28, 45, 30]. These methods quantize images into sequences of discrete tokens (e.g., via VQGAN [44]) and model the token distribution sequentially using cross-entropy loss. While this enables efficient training and fast inference, the quantization process inevitably discards fine-grained image details, limiting reconstruction fidelity. Notable examples include RAR [46], which refines generation by permuting token sequences for richer context, and VAR [34], which models image structure through next-scale prediction. However, the inherent limitations of discrete tokens restrict the expressive capacity needed for high-fidelity synthesis.
@@ 3.3 Proposed Method: DisCon
$\hat{\mathbf{x}}_{d}=\text{DiscreteAR}().$ (6)
@@
Why DisCon Works. We decompose the modeling of continuous tokens into two subproblems. First, we model the discrete tokens $\mathbf{x}_{d}$ using a discrete AR model trained with a cross-entropy loss, a formulation that has been proven effective in prior work. Second, we model the conditional probability from discrete tokens to continuous tokens, i.e., $p(\mathbf{x}_{c}|\mathbf{x}_{d})$, using a continuous AR model. Formally, we factorize the overall distribution as
@@ 4.1 Implementation Details
Table 1: Quantitative comparisons on ImageNet-256. Our method achieves SOTA performance among AR models. Among the metrics, gFID is the most important metric for evaluating the fidelity and diversity of the synthesis result. Params denotes the number of trainable parameters.
@@ 4.3 Ablation Studies
Table 2: Ablation study results on discrete token conditioning. We condition MAR on discrete tokens across various settings, including different parameter sizes and AR steps. Note that these results are preliminary; additional training time and AR steps are expected to further improve performance. The sec/image values were measured using a batch size of 100, and our method’s inference time accounts for the discrete token generation. Params denotes the number of trainable parameters.
@@
Figure 8: Generation Performance during Training. Our method achieves SOTA performance (around 1.5 gFID) after 200 training epochs. Note that these results are obtained with a 0.9999 EMA setting, which may cause a slight performance lag. Detailed training loss curves in the supplementary materials provide additional insights.
@@
Figure 8 shows the generation performance of our method during training. Our model achieves SOTA performance (around 1.5 gFID) after only 200 training epochs, although the use of a 0.9999 EMA (Exponential Moving Average) may introduce a performance lag. This training efficiency demonstrates the reduced optimization complexity achieved by modeling the conditional probability from the discrete distribution to the continuous distribution. Detailed training loss curves in the supplementary materials further support these findings.
@@ 5 Conclusion
Acknowledgments
This work is supported in part by the Hong Kong Research Grants Council under the NSFC/RGC Collaborative Research Scheme (Grant CRS_HKU703/24), the National Natural Science Foundation of China (No. 62202199), and the Fundamental Research Funds for the Central Universities.
Supplementary Materials
Figure 9: Demonstration of Failure Cases. Top: Images generated by RAR-XXL. Bottom: Images generated by our proposed DisCon-L. Although such issues are common in image synthesis, our method exhibits improved performance.
Appendix A Implementation Details
In our experiments, the training is conducted for a default of 800 epochs. For DisCon-L, the batch size per GPU is set to 56, whereas for DisCon-B it is set to 90. Both models share a backbone with 32 transformer blocks and a width of 1024. The primary difference between the two lies in the diffusion head: DisCon-L employs 12 blocks with a width of 1536, while DisCon-B uses 3 blocks with a width of 1024.
Appendix B Failure Cases
Figure 9 illustrates typical failure cases observed in both our method and RAR-XXL, including challenges with human faces, characters, and hands. It is important to emphasize that these issues are inherent to most image generation methods and are not unique to any single approach. Despite the prevalence of these common challenges, our approach consistently outperforms the SOTA RAR-XXL model.
Figure 10: Training Loss Curve of DisCon-L. The loss converges at around 100 epochs, demonstrating the reduced optimization complexity of our approach.
Appendix C Training Process
Figure 10 shows the training loss curve for DisCon-L. The loss stabilizes at around 100 epochs, demonstrating the reduced optimization complexity achieved by our two-stage approach. The efficient training dynamics underscore the benefits of decoupling the modeling of discrete and continuous representations, leading to more reliable and high-quality image synthesis.
Appendix D Discrete AR Models
Our method leverages discrete tokens generated by discrete AR models. We explore performance under different RAR models in Figure 11. Additionally, we present inference results obtained by decoding these discrete tokens, which consistently demonstrate improved performance of our method.
Appendix E Generated Results
Figure 12 presents sample images generated under different class labels, showcasing the high-fidelity synthesis and diversity achieved by our method.
Appendix F Limitations and Future Directions.
Figure 11: Results conditioned on discrete tokens generated by different AR models. From top to bottom: RAR-XXL, RAR-L, and RAR-B. For each model, the top row shows results generated by the respective RAR model, while the bottom row displays outputs from our DisCon method. Zoom in for better visualization to observe the significant improvements in generation quality.
Figure 12: Class-Conditioned Generation. This figure showcases images generated by DisCon-L across various classes, demonstrating the high fidelity and diversity achieved by our approach.